OpenStudy (linn99123):

Evaluate -13 + |3.1 - 4.2|.

OpenStudy (linn99123):

A.) -14.1 B.) -11.9 C.) -5.7

Elsa213 (elsa213):

-13 + |-1.1| -13 + 1.1=-11.9

undeadknight26 (undeadknight26):

I3.1 - 4.2I = I-1.1I = 1.1 + 13 = 14.1

OpenStudy (linn99123):

o.o i really hate numbers now T.T AND LOL

undeadknight26 (undeadknight26):

WAIT! -11.9

Elsa213 (elsa213):

hehehehehe >.>

undeadknight26 (undeadknight26):

i thought it was positive 13 /.\ so -13 + 1.1 = 11.9!

undeadknight26 (undeadknight26):

Unda stand?

OpenStudy (anonymous):

11.9 :/

undeadknight26 (undeadknight26):

^ what he said :D Goo job broda.

OpenStudy (linn99123):

yEA A LITTLE

undeadknight26 (undeadknight26):

-11.1* Oh my goodness i keep forgetting things...I need sleep <(^.O>O.^)>

undeadknight26 (undeadknight26):

-11.9 =.= *kills self*

OpenStudy (linn99123):

go to sleep bruh

undeadknight26 (undeadknight26):

Guest over...and we are playing cards.

Elsa213 (elsa213):

u add c: